SwaggyMark:

Determine the equation of the parabola with focus (-4,5) and directrix x=16.

surjithayer:

parabola is the locus of apoint which is equidistant from focus and directrix. eq. of directrix is x=16 or x-16=0 let P(x,y) be any point on the parabola. Let M be foot of perpendicular from P. focus is F(-4,5) PM=PF \[\frac{ x-16 }{ \sqrt{1^2} }=\sqrt{(x+4)^2+(y-5)^2}\] squaring \[(x-16)^2=(x+4)^2+(y-5)^2\] solve

NineTailedFox:

we just need to find x and y correct? or am i thinking wrong

SwaggyMark:

@ninetailedfox wrote:
we just need to find x and y correct? or am i thinking wrong
we need to find the equation of the parabola given the focus and directrix